Solve the following wave partial differential equation of the vibration of string for ?(? ,?).

yxx=16ytt

y(0,t)=y(1,t)=0

y(x,0)=2sin(\pix)+5sin(3\pix)

yt(x,0)=6\pisin(4\pix)+10\pisin(8\pix)
Answer #1

y (0-4) = y(4.4) Boundary condition Alo antial condition 4(a.o) = avintat 5 sin Bra) Givan that, You = 16 Yff su acău 50 (L1)SO By chifferentiating 46.45 w.vit tinue to givevelocistas y (out) = f (3. –sin Ant (an) + Bo costant.ro] sin mua ft intial vFeel free to use comment section incase you need any help.

Thank you :)